OpenStudy (anonymous):

i need help with this -5x+y=5 find the equations and slop

OpenStudy (slinley):

oh wow, i luvv these problems wait. imma solve it

OpenStudy (anonymous):

First you have to solve for y. So you would move the -5x to the right side of the equation by adding 5x to both sides. Like this -5x+y+5x = 5 +5x. The you have y = 5x+5 and the equation is now in slope intercept form (y=mx+b) therefore the slope of the equation would be 5.